# AI-Driven High-Resolution Cell Segmentation and Quantitative Analysis

### v1

@@ 1 Introduction
A substantial portion of Earth’s microbial biomass—estimated to reach up to 90%—resides in the deep biosphere, particularly in subseafloor sediments under high hydrostatic pressure [1, 2, 3]. While these piezophilic microorganisms dominate high-pressure environments, extremophiles more broadly inhabit a diverse range of ecological niches defined by extremes in temperature, pH, salinity, and radiation [4, 5]. These organisms not only contribute significantly to global biogeochemical cycles but also serve as valuable models for understanding the limits of life in planetary and astrobiological contexts.
Despite their ecological and scientific importance, extremophiles remain poorly characterized at the single-cell level. Several longstanding challenges have hindered their study, including difficulties in cultivation, slow growth rates, and most critically, a lack of reliable genetic tools that prevents the use of traditional reporter-based imaging approaches [6, 7]. As a result, researchers often rely on intrinsic signals such as autofluorescence from metabolic cofactors like NADH and FAD to non-invasively assess cellular activity [8]. In parallel, advances in microscopy now make it possible to acquire high-resolution images of extremophiles in situ, opening new opportunities for quantitative analysis of their physiological states [9].
Among the key parameters accessible through imaging, cell morphology—particularly size and shape—serves as a sensitive indicator of microbial physiology and responses to environmental stress [10, 11]. To extract this information from microscopy data, instance segmentation plays a central role by delineating individual cells with pixel-level precision. However, existing segmentation workflows often rely on manual annotation, which is time-consuming and limits scalability. To address this, there is a growing need for robust, automated segmentation methods that can accommodate the unique challenges of microbial imaging and support high-throughput, quantitative single-cell analysis in extremophiles and beyond.
@@
Despite these advancements, the above-described methods require building domain-specific segmentation datasets with many samples to train a segmentation model, which is expensive and time-consuming for a new cell imaging study. The emergence of segmentation foundation models, such as Segment Anything Model (SAM) [18], enables a paradigm without domain-specific dataset construction or additional training through the zero-shot learning ability. Nevertheless, unique challenges such as noise interference, overlapping cell structures, blurry boundaries, and computational inefficiencies remain significant in achieving accurate segmentation with a foundation model. To this end, we introduce a foundation model-driven cell segmentation pipeline that integrates denoising, segmentation, and post-processing for improved accuracy and quantitative analysis. Our approach enhances image quality using a denoising algorithm to suppress noise while preserving cellular details. It then employs SAM with appropriate prompting for zero-shot instance segmentation without requiring labeled training data. Finally, a post-processing pipeline refines segmentation masks and extracts meaningful quantitative parameters such as cell length, width, and volume. By integrating these techniques, our approach substantially improves segmentation accuracy, minimizes the need for manual intervention, enhances scalability for large-scale biological image analysis, and accelerates the overall processing pipeline. The following sections detail our methodology, present experimental results, and discuss the broader implications of our approach in biomedical research.
Figure 1: Foundation model-driven workflow for extreme imaging. Data Acquisition: A pressure and temperature regulated imaging system that utilizes 2 photon excitation and laser scanning microscopy (a) was used to generate time series images (b). Denoising: These images were averaged to produce a stacked image (c), enhancing signal-to-noise ratio. The stacked image was then denoised using the BM3D algorithm (d). Zero-Shot Segmentation: The denoised image was processed by the Segmentation Anything Model (SAM), a foundation model that generated initial cell masks using point prompts (e). post-processing: post-processing steps were applied to refine the masks (f). Quantitative Analysis: The final masks (g) were aggregated (h) and used for quantitative analysis (i), enabling the extraction of cellular features such as average intensity, cell volume, length, cylindrical height, and average width (j).
@@ 2 Methodology
Fig. 1 shows the AI-driven workflow for automated cell segmentation and quantitative morphological analysis under extreme imaging conditions. By running this AI-driven cell imaging system, we can automatically obtain detailed cellular features that are essential for studying microbial growth and metabolism, providing critical insights into their evolutionary adaptations to harsh environments. Detailed descriptions of each component of the workflow are provided in the following subsections.
@@ 2.1 Data Acquisition
To minimize photobleaching and phototoxicity—critical factors in quantitative fluorescence microscope techniques such as scanning number and brightness (sN&B) and fluorescence lifetime imaging microscopy (FLIM), a pixel dwell time of 40 $\mu$s was used. A total of seven raster scans were collected per field of view (FOV), where the number of scans can be adjusted depending on the fluorescence imaging technique used. Each FOV comprised 256 $\times$ 256 pixels, corresponding to a physical area of 20 $\mu$m $\times$ 20 $\mu$m. This scale accommodates approximately 30 yeast cells or 100–200 bacterial cells per FOV, depending on cell size, shape, and species. Although imaging under extreme conditions such as high pressure presents additional challenges [19], the imaging process relies on the same software and hardware, making the AI-driven segmentation workflow applicable not only to rod-shaped extremophiles but broadly to all rod-shaped microbes.
@@ 2.3.1 Segment Anything Model (SAM)
SAM is a foundation model designed for general-purpose segmentation. It operates in a prompt-based manner, where segmentation masks are generated based on user-provided prompts such as points, bounding boxes, or free-form text descriptions. The model consists of three primary components: an image encoder, a prompt encoder, and a mask decoder. The image encoder, a vision transformer (ViT), takes a set of none-overlap patches as inputs and outputs their feature vectors to represent the input image. The prompt encoder processes user-defined prompts, which can be masks, points, boxes, or text, and outputs the prompt embedding vectors. Given the extracted image features and encoded prompts, the mask decoder will predict the segmentation masks and Intersection-over-Union (IoU) scores.
@@
A key advantage of SAM is its zero-shot learning capability, enabling it to segment objects in previously unseen datasets without requiring task-specific fine-tuning. By leveraging large-scale pretraining, SAM generalizes well across a wide range of domains—including natural images and biomedical microscopy data—demonstrating strong versatility across different visual contexts [18].
@@ 2.3.2 Enhanced Segment Anything Model (SAM2)
Post-processing is an essential step to refine segmentation results by eliminating artifacts, reducing redundancy, and enhancing accuracy. The raw segmentation outputs from SAM and SAM2 may contain small artifacts, duplicate masks, partial segmentations, or edge-touching objects, requiring additional processing before quantitative analysis. To address these issues, we apply a structured post-processing pipeline that filters, refines, and optimizes the segmented masks. The post-processing workflow consists of the following steps.
Initial Mask Filtering Based on Area and Intensity: We set thresholds for area and intensity to remove small artifacts and irrelevant objects. Masks smaller than 100 pixels are discarded as noise, while those exceeding 1250 pixels are removed as they may correspond to merged objects or background artifacts. Additionally, we apply intensity filtering to eliminate artifacts caused by incorrect segmentation. The mean intensity of all selected cells ($\bar{I}$) is computed, and masks deviating significantly from this mean, defined as $0.35\bar{I}<I_{i}<1.6\bar{I}$, are removed.
Removing Fully Contained Masks: Masks that are entirely enclosed within larger ones are redundant and can lead to overestimation of object counts. Given two masks $M_{i}$ and $M_{j}$, if $M_{i}\subseteq M_{j}$, the smaller mask is discarded. This prevents duplicate segmentations of the same object from being counted multiple times.
Non-Maximum Suppression (NMS) for Overlapping Masks: Highly overlapping masks may result in over-segmentation. To address this, we apply Non-Maximum Suppression (NMS) by computing the Intersection-over-Union (IoU) between each pair of masks. If two masks $M_{i},M_{j}$ have a high IoU, given by
$\text{IoU}(M_{i},M_{j})=\frac{|M_{i}\cap M_{j}|}{|M_{i}\cup M_{j}|},$ (1)
then the smaller mask is removed to ensure that only one high-confidence segmentation is retained per object. In this study, the IoU threshold is set as 0.3.
@@
where $\oplus$ represents dilation and $\ominus$ represents erosion with structuring element $K$. This operation ensures a more precise cell shape representation.
By integrating this post-processing pipeline, we achieve more reliable segmentation results, better-defined cell boundaries, reduced redundant masks, and improved accuracy in quantitative cell analysis.
@@ 3.2 Segmentation Results
The application of BM3D denoising leads to a substantial improvement in segmentation quality, particularly for SAM-L and SAM-B. Enhanced contrast and reduced background noise allow these models to better resolve cell boundaries and produce a greater number of valid segmentation masks. These observations underscore the importance of input quality: with sufficient denoising, all SAM variants demonstrate notable performance improvements, with SAM-L maintaining a modest advantage in segmentation accuracy.
Across all tested conditions, SAM-H consistently yields the most accurate and complete segmentation. In contrast, SAM-B and SAM-L frequently fail to capture finer morphological details, as evidenced by the white ellipses in Figs. 3(e) and (h). However, SAM-H is not without limitations. As shown in Fig. 4(b), it occasionally exhibits over-segmentation—splitting a single cell into multiple masks—as illustrated by the white-circled examples, where the closely adjacent cells are both segmented as two overlapped objects.
@@ 4 Quantitative Features
Fig. 7 illustrates the quantitative feature extraction for a single cell. The average intensity and length were computed by overlaying the cell mask onto the stacked image, as shown in Fig. 7(a), resulting in an average intensity of 2.698 and a length of 3.215 µm. The average width was measured from the central two rectangles, and the cell volume was estimated based on a geometric model consisting of a cylinder and two hemispheres. As shown in Fig. 7(b), the average width was 0.865 µm, the cylinder height (length excluding the hemispherical ends) was 2.350 µm, and the calculated cell volume was 1.720 fL.
Figure 7: Illustration of quantitative feature extraction for a single cell. (a) The cell length was computed from the major axis of the red bounding box surrounding the cell, while the average intensity was calculated by overlaying the cell mask onto the stacked image, resulting in a length of 3.215 µm and an average intensity of 2.698. (b) The average width was estimated from the central two bounding rectangles, and the cell volume was computed using a geometric model consisting of a cylinder and two hemispherical caps. The derived features included an average width of 0.865 µm, a cylinder height of 2.350 µm, and a cell volume of 1.720 fL.
@@ 5 Conclusion
In this study, we introduced an AI-driven pipeline for high-resolution microscopy image segmentation and analysis, incorporating BM3D-based denoising, the Segment Anything Model (SAM), a structured post-processing routine, and a quantitative analysis workflow. This integrated approach effectively addresses key challenges in biological imaging, including noise, low contrast, and overlapping cell structures. Quantitative evaluations show that the SAM-Huge variant, when paired with denoising and post-processing, achieves the highest segmentation accuracy, with an average error rate of just 3.0%. In contrast, SAM2-Large performs less effectively in resolving fine structural details under low-contrast conditions. Beyond segmentation, our pipeline enables precise extraction of essential morphological features—including intensity, length, width, and volume—through a geometric modeling approach. These capabilities facilitate in-depth studies of microbial metabolism, structural adaptation, and evolutionary dynamics in extreme environments. Overall, the proposed framework provides a scalable, label-free, and high-fidelity solution for quantitative cellular analysis, with strong potential for extension to a wide range of biomedical imaging applications requiring accurate structural characterization.

### v2

# AI-Driven Segmentation and Analysis of Microbial Cells

@@ 1 Introduction
A substantial portion of Earth’s microbial biomass resides in the extreme conditions such as high pressure, temperature, salinity, or pH [1, 2, 3, 4, 5]. They are key players in global biogeochemical cycles and serve as models for understanding the limits of life in planetary and astrobiological contexts[4], and offer insight into microbial evolution under high-pressure, early Earth–like conditions[6]. Yet, their characterization at the single-cell level remains limited due to difficulties in cultivation, slow growth, and the lack of genetic tools for conventional reporter-based imaging [7, 8]. To overcome these barriers, researchers increasingly rely on intrinsic metabolic autofluorescence (e.g., NADH, FAD, F420) and recent advances in microscopy also enable high-resolution, in situ imaging and quantitative analysis of cells in their native environments [9, 10].
Among the key parameters accessible through imaging, cell morphology—particularly size and shape—serves as a sensitive indicator of microbial physiology and responses to environmental stress [11, 12, 13]. To extract this information from microscopy data, instance segmentation plays a central role by delineating individual cells with pixel-level precision. However, existing segmentation workflows often rely on manual annotation, which is time-consuming and limits scalability. To address this, there is a growing need for robust, automated segmentation methods that can accommodate the unique challenges of microbial imaging and support high-throughput, quantitative cell analysis in extremophiles and beyond.
@@
Despite these advancements, the above-described methods require building domain-specific segmentation datasets with many samples to train a segmentation model, which is expensive and time-consuming for a new cell imaging study. The emergence of segmentation foundation models, such as the Segment Anything Model (SAM) [20], enables a paradigm without domain-specific dataset construction or additional training through the zero-shot learning ability. Nevertheless, unique challenges such as noise interference, overlapping cell structures, blurry boundaries, and computational inefficiencies remain obstacles to achieving accurate segmentation with a foundation model. To this end, we introduce a foundation model-driven cell segmentation pipeline that integrates denoising, segmentation, and post-processing for fast, accurate and quantitative analysis. Our approach enhances image quality using a denoising algorithm to suppress noise while preserving cellular details. It then employs SAM with appropriate prompting for zero-shot instance segmentation without requiring labeled training data. Finally, a post-processing pipeline refines segmentation masks and extracts meaningful quantitative parameters such as cell length, width, and volume. By integrating these techniques, our approach substantially improves segmentation accuracy, minimizes the need for manual intervention, enhances scalability for large-scale biological image analysis, and accelerates the overall processing pipeline. The following sections detail our methodology, present experimental results, and discuss the broader implications of our approach in biomedical research.
Figure 1: Foundation model-driven workflow for cell imaging. Data Acquisition: A pressure and temperature regulated imaging system that utilizes 2 photon excitation and laser scanning microscopy (a) was used to generate time series images (b). Denoising: These images were averaged to produce a stacked image (c), enhancing signal-to-noise ratio. The stacked image was then denoised using the BM3D algorithm (d). Zero-Shot Segmentation: The denoised image was processed by the Segmentation Anything Model (SAM), a foundation model that generated initial cell masks using point prompts (e). Post-processing: Post-processing steps were applied to refine the masks (f). Quantitative Analysis: The final masks (g) were aggregated (h) and used for quantitative analysis (i), enabling the extraction of cellular features such as average intensity, length, average width, and cell volume (j).
@@ 2 Methodology
Fig. 1 shows the AI-driven workflow for automated cell segmentation and quantitative morphological analysis. By running this AI-driven cell imaging system, we can automatically obtain detailed cellular features that are essential for studying microbial growth and metabolism, providing critical insights into their evolutionary adaptations to harsh environments. Detailed descriptions of each component of the workflow are provided in the following subsections.
@@ 2.1 Data Acquisition
To minimize photobleaching and phototoxicity—critical factors in quantitative fluorescence microscope techniques such as scanning number and brightness (sN&B) and fluorescence lifetime imaging microscopy (FLIM), a pixel dwell time of 40 $\mu$s was used. A total of seven raster scans were collected per field of view (FOV), where the number of scans can be adjusted depending on the fluorescence imaging technique used. Each FOV comprised 256 $\times$ 256 pixels, corresponding to a physical area of 20 $\mu$m $\times$ 20 $\mu$m. This scale accommodates approximately 30 yeast cells or 100–300 bacterial cells per FOV, depending on cell size, shape, and species. Although imaging under extreme conditions such as high pressure presents additional challenges [10], the acquired data will always have the same file type (i.e. TIFF), making the AI-driven segmentation workflow applicable irrespective of hardware, software, or imaging conditions.
@@ 2.3.1 Segment Anything Model (SAM)
SAM is a foundation model designed for general-purpose segmentation. A key advantage of SAM is its zero-shot learning capability, enabling it to segment objects in previously unseen datasets without requiring task-specific fine-tuning. By leveraging large-scale pretraining, SAM generalizes well across a wide range of domains—including natural images and biomedical microscopy data—demonstrating strong versatility across different visual contexts [20].
SAM operates in a prompt-based manner, where segmentation masks are generated based on user-provided prompts such as points, bounding boxes, or free-form text descriptions. The model consists of three primary components: an image encoder, a prompt encoder, and a mask decoder. The image encoder, a vision transformer (ViT), takes a set of none-overlap patches as inputs and outputs their feature vectors to represent the input image. The prompt encoder processes user-defined prompts, which can be masks, points, boxes, or text, and outputs the prompt embedding vectors. Given the extracted image features and encoded prompts, the mask decoder will predict the segmentation masks and Intersection-over-Union (IoU) scores.
@@ 2.3.2 Enhanced Segment Anything Model (SAM2)
Post-processing is an essential step to refine segmentation results by eliminating artifacts, reducing redundancy, and enhancing accuracy. The raw segmentation outputs from SAM and SAM2 may contain small artifacts, duplicate masks, partial segmentations, or edge-touching objects, requiring additional processing before quantitative analysis. To address these issues, we apply a structured post-processing pipeline that filters, refines, and optimizes the segmented masks. The post-processing workflow consists of the following steps:
Initial Mask Filtering Based on Area and Intensity: We set thresholds for area and intensity to remove small artifacts and irrelevant objects. Masks smaller than 100 pixels are discarded as noise, while those exceeding 1250 pixels are removed as they may correspond to merged objects or background artifacts. Additionally, we apply intensity filtering to eliminate artifacts caused by incorrect segmentation. The mean intensity of all selected cells ($\bar{I}$) is computed, and masks deviating significantly from this mean, defined as $0.35\bar{I}<I_{i}<1.6\bar{I}$, are removed. These thresholds are selected through preliminary experimentation.
Removing Fully Contained Masks:When using SAM, which tends to over-segment objects by treating distinct parts as separate instances. For example, in natural images, SAM may segment a car into multiple overlapping regions such as the body, windows, and tires. Similarly, in microscopy images, a single cell can be incorrectly segmented into one large mask along with one or two smaller interior masks. To mitigate this, we discard smaller masks that are completely contained within larger ones. Formally, given two masks $M_{i}$ and $M_{j}$, if $M_{i}\subseteq M_{j}$, the smaller mask is removed to prevent duplicate segmentations of the same object from being counted multiple times.
Non-Maximum Suppression (NMS) for Overlapping Masks: To address over-segmentation caused by closely overlapping masks, we implement Non-Maximum Suppression (NMS) based on pairwise Intersection over Union (IoU). For each mask $M_{i}$, we compute the IoU with all other masks $M_{j}$, defined as:
$\text{IoU}(M_{i},M_{j})=\frac{|M_{i}\cap M_{j}|}{|M_{i}\cup M_{j}|}.$ (1)
If the IoU between $M_{i}$ and $M_{j}$ exceeds a threshold (set to 0.3 in this study), the mask with the smaller area is removed. This approach ensures that among highly overlapping masks, only the largest and most representative one is retained, thereby preventing redundant segmentation of the same object.
@@
where $\oplus$ represents dilation and $\ominus$ represents erosion with structuring element $K$. This operation ensures a more precise cell shape representation. By integrating this post-processing pipeline, we achieve more reliable segmentation results, better-defined cell boundaries, reduced redundant masks, and improved accuracy in quantitative cell analysis.
@@ 3.2 Segmentation Results
The application of BM3D denoising leads to a substantial improvement in segmentation quality, particularly for SAM-L and SAM-B. Enhanced contrast and reduced background noise allow these models to better resolve cell boundaries and produce a greater number of valid segmentation masks. These observations underscore the importance of input quality: with sufficient denoising, all SAM variants demonstrate notable performance improvements, with SAM-L maintaining a modest advantage in segmentation accuracy. Across all tested conditions, SAM-H consistently yields the most accurate and complete segmentation. In contrast, SAM-B and SAM-L frequently fail to capture finer morphological details, as evidenced by the white ellipses in Figs. 3(e) and (h). However, SAM-H is not without limitations. As shown in Fig. 4(b), it occasionally exhibits over-segmentation—splitting a single cell into multiple masks—as illustrated by the white-circled examples, where the closely adjacent cells are both segmented as two overlapped objects.
@@ 4 Quantitative Features
Fig. 7 illustrates the quantitative feature extraction for a single cell. The average intensity and length were computed by overlaying the cell mask onto the stacked image, as shown in Fig. 7(a), resulting in an average intensity of 2.698 and a length of 3.215 µm. The average width was measured from the central two rectangles, and the cell volume was estimated based on a geometric model consisting of a cylinder and two hemispheres. As shown in Fig. 7(b), the average width was 0.865 µm,and the calculated cell volume was 1.720 fL.
Figure 7: Illustration of quantitative feature extraction for a single cell. (a) The cell length was computed from the major axis of the red bounding box surrounding the cell, while the average intensity was calculated by overlaying the cell mask onto the stacked image, resulting in a length of 3.215 µm and an average intensity of 2.698. (b) The average width 0.865 µm was estimated from the central two bounding rectangles, and the cell volume 1.720 fL was computed using a geometric model consisting of a cylinder and two hemispherical caps.
@@ 5 Conclusion
In this study, we introduced an AI-driven pipeline for high-resolution microscopy image segmentation and analysis, incorporating BM3D-based denoising, the Segment Anything Model (SAM), a structured post-processing routine, and a quantitative analysis workflow. This integrated approach effectively addresses key challenges in biological imaging, including noise, low contrast, and overlapping cell structures. Quantitative evaluations show that the SAM-Huge variant, when paired with denoising and post-processing, achieves the highest segmentation accuracy, with an average error rate of just 3.0%. In contrast, SAM2-Large performs less effectively in resolving fine structural details under low-contrast conditions. Beyond segmentation, our pipeline enables precise extraction of essential morphological features—including intensity, length, width, and volume—through a geometric modeling approach. These capabilities would facilitate in-depth studies of microbial metabolism, structural adaptation, and evolutionary dynamics in extreme environments. Overall, the proposed framework provides a scalable, label-free, and high-fidelity solution for quantitative cellular analysis, with strong potential for extension to a wide range of biomedical imaging applications requiring accurate structural characterization.
@@ Acknowledgments
This research is supported by the DOE Office of Science, Office of Biological and Environmental Research (BER), grant no. 280597.